Exercise 24-4 Partially correct answer. Your answer is partially correct. Try again. BAK Corp. is considering purchasing one of two new diagnostic machines. Either machine would make it possible for the company to bid on jobs that it currently isn’t equipped to do. Estimates regarding each machine are provided below. Machine A Machine B Original cost $76,900 $186,000 Estimated life 8 years 8 years Salvage value 0 0 Estimated annual cash inflows $19,600 $39,800 Estimated annual cash outflows $5,150 $10,080 Click here to view PV table. Calculate the net present value and profitability index of each machine. Assume a 9% discount rate. (If the net present value is negative, use either a negative sign preceding the number eg -45 or parentheses eg (45). Round answer for present value to 0 decimal places, e.g. 125 and profitability index to 2 decimal places, e.g. 10.50. For calculation purposes, use 5 decimal places as displayed in the factor table provided.) Machine A Machine B Net present value Entry field with incorrect answer Entry field with incorrect answer Profitability index Entry field with incorrect answer Entry field with incorrect answer Which machine should be purchased?
Answer #1

Net annual cash inflow = Cash inflow - Cash outflow

Machine A = $19,600-5,150 = $14,450

Machine B = $39,800-10,080 = $29,720

Net present value = Present value of cash inflow - initial investment

Machine A = [$14,450*5.53482 (PVF@9%,8 years)] $79,978 - 76,900 = $3,078

Machine B = $29,720*5.53482 = $164,495 - 186,000 = -$21,505

Profitability index = Present value of cash flow / initial investment

Machine A = $79,978 / 76,900 = 1.04

Machine B = $164,495 / 186,000 = 0.88

As per above results, Machine A should be purchased.